OpenStudy (anonymous):

integrate:sin(y),y,0,(1/2)

OpenStudy (anonymous):

intergral of sin(y) = -cos(y)

OpenStudy (anonymous):

-cos(1/2)-(-cos(0))

OpenStudy (anonymous):

-cos(1/2)+cos(0)

OpenStudy (anonymous):

with respect to what? is its respect to y then integral of sin(y) = -cos(y) integral of y = 1/2 y^2 integral of 0 = C integral of 1/2 = 1/2x

OpenStudy (anonymous):

I first had to integrate arcsin(x) with respect to x from 0 to 1/2 using integration by parts....then it ask me to Calculate the area a second way by integrating with respect to y (using horizontal rectangles to approximate area instead of vertical rectangles.)

OpenStudy (anonymous):

was i right to assume this ment i noew integrate sin(y)=x with respect to y from 0 to 1/2?

OpenStudy (anonymous):

that should be right

OpenStudy (anonymous):

what is right?

OpenStudy (anonymous):

to do it another way yes you can use horizontal rectsangle and integrate with respect to y and it will give you the same answer

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

thank you though